OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of two-thirds?

OpenStudy (anonymous):

I presume by standard from you me y=mx+b so i need slope (which you have and b (y intercept) so we have x and y in (6, 3) and m in slope 2/3 so plug it in 3=(2/3)6+b and solve for b 3= 12/3 +b 3=4 +b -1 = b now write it in the form \[y= \frac{ 2 }{ 3 }x -1\]

OpenStudy (anonymous):

The form you used is slope-intercept form. The standard form is Ax+By=C, where A, B, and C are all integers.

OpenStudy (anonymous):

im not good at algebra lol. ill try, thanks for the help